The triangle ABC is right-angled with a right angle at C. The height CD has a length of 5 units of length (D lies on the side AB), and the distance AD ​​has a length of 7 units of length. Determine and state the length of the hypotenuse. Briefly explain each step


1
Expert's answer
2021-05-07T11:44:43-0400

According to given information,

we can draw diagram as follows:

with some variables x, b and p.





as CD is height so CD is perpendicular to AB.


NOW,

applying Pythagoras theorem in triangle ADC,


"7^2+5^2=p^2"


"p=\sqrt{74}" ............................(1)


applying Pythagoras theorem in triangle BDC,



"x^2+5^2=b^2" ...............(2)


applying Pythagoras theorem in triangle ACB ,using (1),


"74+b^2=(7+x)^2" ...........................(3)



eliminating b from eqn. (2) and(3)...

we get,


"x^2 +25=(7+x)^2-74"


"\boxed {x={50\over 14}}"


therefore,

length of hypotenuse is,


="x+7"

"={50\over 14} +7"


"\boxed{ = 10.57 \ unit}"